OpenStudy (anonymous):

find the exact area between y=ln(x) and y=ln(x^2) for 1<= x <= 2

OpenStudy (anonymous):

i will guess \[\int_1^22\ln(x)-\ln(x)dx=\int_1^2\ln(x)dx\]

OpenStudy (anonymous):

anti derivative is \[x\ln(x)-x\] evaluate at 2 and 1 and subtract